OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. −6x + 3y = 18 4x − 3y = 6 2x + 4y = 24 6x + 12y = 36 3x − y = 14 −9x + 3y = −42 5x + 2y = 13 −x + 4y = −6

OpenStudy (anonymous):

@campbell_st @CBARREDO1

OpenStudy (campbell_st):

well look at the pairs of equations and determine if 1 equation is a multiple of the other equation... wehn you have a multiple line x + y = 2 and -2x -2y = -4 you have the same equation, hence there is an infinite number of solutions... one method might be to graph each pair... and see which pair only produce 1 line this is a good site for graphing https://www.desmos.com/calculator hope it helps

OpenStudy (cbarredo1):

3x-y=14, because 3x-y= 14 multiplied by -3 is -9x+3y= -42 so the equations are equal